DIRC Operations Weekly Report
Period: Wednesday 2007/6/13 - Wednesday 2007/6/19
Oncall: Alejandro Perez
Fast Monitoring: Justine Serrano, Alejandro Perez
Luminosities this week
PEP peak lumi: ~1 x 1034
(last week: ~9.5 x 1033)
Run 6 PEP delivery: 2.27 fb-1 Offpeak
(last weeks: 4.8 fb-1)
BaBar integrated lumi:2.2 fb-1 Offpeak
(last weeks: 4.6 fb-1)
Run 6 total delivery: 49.10 fb-1
Run 6 total BaBar recorded: 47.29 fb-1 total, including 4.18 fb-1 off-peak data
PEP activities/issues this week
Hot weather has disrupted datatime operations, limited data integration.
PEP had many HER aborts. To be addressed on 06/21 ROD.
06/15: Magnet PS tripped around 6.30am. Recoverred by 8am.
06/19: After taking 4.18 fb-1 of offpeak data PEP switched to Onpeak running.
BaBar activities/issues this week
Some subsystems have raised temperature thresholds because of hot weather.
06/15: HVAC rebooted their control computer and all the central chiller went off (accidentally).
DIRC activities/issues this week
06/14: A flaky temperature sensors (in sector 10) started to give many yellow alarms due to heat wave. We raised the inlet temperature yellow alarm threshold for this sector (from 34 to 35 degrees).
06/15: Due to central chiller problem the DIRC chiller tripped and FEE went off. DIRC recovered fine after HVAC people turned central chillers back on.
06/19: 8 new channels masked out (the first 8 channels of DFB 0, sector 0).
See DIRC logbook for details
Lumi and Background this week
- ROD scheduled for June 21 (~2 shifts):
- We will remove the scope monitoring chiller power.
- Ray will replace a board in the back of DIRC SIAM.
- Replace six DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 6, DFB 1.
- sector 8, DFB 0.
- X-tree in sector 7 HV group 55 that only shows up during cosmics.
- X-tree in sector 0 HV group 1.
Alejandro and Justine rotating backup & on-call commissioning duties.